OpenStudy (anonymous):

find the equation of a sphere whose centre is (2,3,1) and touches the plane x+y+z=0

OpenStudy (anonymous):

Can you compute the distance between the plane and the centre of the sphere? HNF: \[\Large \frac{x+y+z}{\sqrt{3}}=0 \]